OpenStudy (anonymous):

Pleeaasse help! Write an equation for the line through (0, 0) and (2, 2) in standard form.

OpenStudy (anonymous):

k=(y2-y1)/(x2-x1)= y=k(x-x1)+y1=

OpenStudy (loser66):

Actually, you have to find the slope, and then put the point and the slope into the formula of the equation of the line. However, You have an advantage on those points and it give you immediately the equation of the line. You can see (0,0) 0=0, right? and (2,2) 2 =2 ,right? so that the equation is y =x done